# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v1

@@ Abstract
We study continuous-time portfolio choice in diffusion markets with parameter $\theta\in\Theta$ and uncertainty law $q(d\theta)$. Nature draws latent $\theta\sim q$ at time 0; the investor cannot observe it and must deploy a single $\theta$-blind feedback policy maximizing an ex–ante CRRA objective averaged over diffusion noise and $\theta$. Our methods access $q$ only by sampling and assume no parametric form. We extend Pontryagin–Guided Direct Policy Optimization (PG–DPO) by sampling $\theta$ inside the simulator and computing discrete-time gradients via backpropagation through time (BPTT), and we propose projected PG–DPO (P–PGDPO) that projects costate estimates to satisfy the $q$-aggregated Pontryagin first-order condition, yielding a deployable rule. We prove a BPTT–PMP correspondence uniform on compacts and a residual-based $\theta$-blind policy-gap bound under local stability with explicit discretization/Monte Carlo errors; experiments show projection-driven stability and accurate decision-time benchmark recovery in high dimensions.
@@ 1 Introduction
A central problem in quantitative finance is to allocate wealth dynamically across many risky assets in continuous time. In the classical Merton model, investment opportunities are described by a low-dimensional diffusion with known drift and volatility, and the investor solves a Hamilton–Jacobi–Bellman (HJB) equation to obtain closed-form optimal portfolios and value functions; see, for example, merton1969lifetime; merton1971optimum and the subsequent literature. In realistic markets, however, neither expected returns nor volatilities are known: they must be estimated from finite samples, often using many assets and predictors and under nontrivial model selection and regularization. Empirically, expected-return forecasting is fragile and return predictability is unstable across samples and over time, with many proposed predictors delivering limited out-of-sample gains (e.g., goyalwelch2008; campbellthompson2008; rapach2010forecast; danglhalling2012tvp; lettauvanNieuwerburgh2008shifts; pettenuzzo2014constraints). In such settings it is crucial to distinguish diffusion risk (Brownian noise conditional on parameters) from statistical parameter uncertainty (error in estimated coefficients). A long line of portfolio-choice work shows that return predictability, learning, and parameter uncertainty can induce substantial intertemporal hedging effects and more conservative allocations (e.g., kandelstambaugh1996predictability; barberis2000investor; campbell2002strategic; brandt2005portfolio; brennan1998investor; xia2001learning; maenhout2004robust).
@@
We study continuous-time portfolio choice when market dynamics are known only up to an estimated parameter $\theta\in\Theta$, where the estimation pipeline produces a nondegenerate uncertainty law $q(d\theta)$ over $\Theta$. We treat $q$ as an input object that encapsulates all statistical information available at time 0: it may be derived from resampling approximations (e.g. bootstrap) (e.g., efron1979bootstrap; efron1994bootstrap), model averaging or Bayesian model uncertainty pipelines (e.g., pastor2000portfolio; avramov2002stock; cremers2002stock), approximate posteriors, or other procedures. Our goal is not to revisit inference, but to optimize decisions given this uncertainty description. Algorithmically, we interact with $q$ only through sampling $\theta\sim q$ inside the simulator; we do not assume closed-form densities, conjugate updates, or any particular parametric form. Concretely, we seek portfolio policies that maximize terminal CRRA utility ex–ante, averaging over both diffusion noise and parameter draws $\theta\sim q$. This formulation also supports offline diagnostics that quantify how recommended allocations vary across the statistically plausible models encoded in $q$.
@@
A key modeling choice is that $\theta$ is latent: Nature draws a fixed but unobserved $\theta\sim q$ at time 0 (independent of the Brownian drivers) and keeps it fixed on $[0,T]$. While the investor knows $q$, she does not observe the realized $\theta$ and must therefore deploy a single $\theta$-blind policy. We restrict attention to Markov feedback policies of the form $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ that depend only on observable wealth $X_{t}$ and market factors $Y_{t}$, and we do not augment the state by a belief/posterior process. This fixed-$q$ commitment is intended as a decision-time benchmark: it targets a single deployable rule given an exogenous uncertainty law, cleanly decoupling how uncertainty is produced (any pipeline yielding $q$) from how decisions are optimized (our solver given $q$). We do not claim that belief-state control is conceptually inappropriate; rather, it defines a different (and typically far more demanding) problem than computing a single $\theta$-blind deployable feedback rule from a fixed uncertainty description (e.g., bensoussan1985optimal; pham2017dynamic). At the same time, fixed-$q$ optimization couples heterogeneous market models: gradient signals can vary substantially across $\theta$ draws and may partially cancel when learning a single global policy end-to-end.
@@
The $\theta$-blind constraint also changes what a first-order optimality condition means. If $\theta$ were observable, Pontryagin’s Maximum Principle (PMP) yields a $\theta$-conditional criticality condition and an associated $\theta$-conditional full-information feedback map (infeasible under latent $\theta$). Under $\theta$-blind deployability, admissible perturbations are also $\theta$-blind. Taking the first variation of the ex–ante objective and using Fubini’s theorem shows that the correct necessary condition is $q$-aggregated: the expectation over $\theta\sim q$ of the Hamiltonian gradient $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ must vanish along the state process, in the standard stochastic maximum principle framework (e.g., yong1999stochastic; fleming2006controlled; pham2009continuous). Because $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ is affine in $\pi$ for our portfolio Hamiltonian, this aggregation yields a statewise linear system whose solution defines a deployable $\theta$-blind projected portfolio rule. Notably, the condition and resulting projection are agnostic to the internal construction of $q$ and depend only on its role as the ex–ante mixing law.
These features place the problem outside the practical reach of classical dynamic programming in the high-dimensional regime we target. In low-dimensional deterministic-parameter Markov models, DP/HJB is canonical; however, even with several factors it requires solving an HJB equation in the state $(t,X_{t},Y_{t})$, where grid-based PDE methods are quickly defeated by the curse of dimensionality (e.g., bellman1961adaptive; kushner2001numerical). Deep PDE surrogates such as PINNs (e.g., raissi2019physics; sirignano2018dgm) and deep BSDE methods (e.g., han2018solving; beck2019machine) alleviate the need for grids, but fully nonlinear portfolio HJBs with many assets and factors remain numerically delicate, especially when accurate mixed derivatives are required. If one further models parameter uncertainty via belief-state augmentation, the state becomes a posterior measure and the control problem becomes infinite-dimensional (e.g., bensoussan1985optimal; pham2017dynamic).
@@
Our approach is simulation-based and builds on Pontryagin–Guided Direct Policy Optimization (PG–DPO) (huh2025breaking; huh2025constraint). PG–DPO parameterizes a $\theta$-blind feedback policy via a neural network, simulates trajectories of the controlled SDE, and employs backpropagation through time (BPTT) to compute exact gradients of terminal utility. Crucially, intermediate pathwise sensitivities computed by BPTT coincide with the stochastic costates (adjoints) in PMP, mirroring the classical duality between backpropagation and adjoint methods (see lecun1988optimal; yong1999stochastic). In the latent-parameter setting, we approximate the ex–ante objective by sampling $\theta\sim q$ inside the simulator and fixing it along each trajectory, while the policy depends only on observable states. To stabilize learning under heterogeneous $\theta$ draws, we extend the projected variant, P–PGDPO, to latent $\theta$: after a warm-up phase that stabilizes costate estimates, we project Monte Carlo Pontryagin objects onto the $q$-aggregated Pontryagin first-order condition. This reconstruction yields a robust deployable $\theta$-blind rule obtained from the $q$-aggregated criticality, and can be amortized into a fast-to-evaluate policy.
In high-dimensional scaling experiments under static Gaussian drift uncertainty, the two-stage projected pipeline substantially improves decision-time accuracy relative to end-to-end learning, with clear stabilization effects in aligned regimes. In misaligned regimes, projection gains diminish with dimension; diagnostics indicate that deterioration is driven primarily by growth of aggregated first-order residuals and curvature mismatch rather than by catastrophic numerical inversion. In factor-driven markets with mean-reverting investment opportunities where return–factor correlation induces intertemporal hedging demand, the projected pipeline recovers the analytic decision-time benchmark under the same $\theta$-blind deployability restriction, while a model-free PPO baseline remains far from the reference in the regimes we test.
Our main theoretical guarantee is a residual-based ex–ante $\theta$-blind policy-gap bound for the deployable fixed-$q$ commitment problem: under mild slab-wise local stability conditions for the $q$-aggregated projection map, a small warm-up aggregated first-order residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal interior deployable $\theta$-blind policy, up to discretization and Monte Carlo error.
@@
Our contributions are threefold. (i) We formulate a latent-parameter, fixed-$q$ ex–ante CRRA portfolio problem under a deployable $\theta$-blind Markov feedback restriction and derive the corresponding $q$-aggregated Pontryagin first-order condition, emphasizing an inference-agnostic interface where uncertainty enters only through an exogenous mixing law $q(d\theta)$. (ii) We extend PG–DPO to this setting by sampling $\theta$ only inside the simulator and using BPTT to compute exact discrete-time gradients and pathwise sensitivities, and we establish a conditional BPTT–PMP correspondence uniform over $\theta$ on compact subsets of $\Theta$. (iii) We develop uncertainty-aware P–PGDPO that projects Monte Carlo costate estimates to produce a deployable $q$-aggregated $\theta$-blind rule, together with a residual-based ex–ante $\theta$-blind policy-gap bound and empirical evidence of two-time-scale stabilization and stability gains from projection.
@@
The remainder of the paper is organized as follows. Section 2 formulates the fixed-$q$ ex–ante portfolio problem under a latent parameter and a deployable $\theta$-blind Markov feedback restriction, and derives the $\theta$-conditional versus $q$-aggregated Pontryagin first-order conditions together with Gaussian decision-time reference models. Section 3 develops PG–DPO and uncertainty-aware P–PGDPO for the latent-$\theta$ setting, establishes the conditional BPTT–PMP correspondence, and proves a residual-based ex–ante $\theta$-blind policy-gap bound under local stability of the aggregated projection map. Section 4 reports high-dimensional scaling experiments under static Gaussian drift uncertainty, and Section 5 studies hedging-demand recovery in factor-driven markets with mean-reverting investment opportunities. Technical proofs and implementation details are collected in the appendix.
@@ 2 Dynamic Portfolio Choice in Estimated Diffusion Markets with Latent Parameter Uncertainty
In this section we formulate a continuous-time dynamic portfolio choice problem with CRRA preferences in a diffusion market whose coefficients are estimated from data and are therefore statistically uncertain. Rather than committing to a particular estimation architecture, we treat the market as belonging to a (possibly high-dimensional) parameterized family indexed by $\theta\in\Theta$, and we represent the uncertainty in the estimated parameter by an exogenously given probability law $q(d\theta)$ over $\Theta$.
  • Nature draws a fixed but unobserved $\theta\sim q$ at time 0 and keeps it constant on $[0,T]$.
  • The investor knows $q$ but does not observe the realized $\theta$ and must deploy a single $\theta$-blind portfolio policy.
  • Performance is evaluated ex–ante by averaging terminal utility over both diffusion noise and $\theta\sim q$.
  • We restrict to $\theta$-blind Markov feedback rules $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ and do not augment the state by a belief/posterior process.
This fixed-$q$, $\theta$-blind formulation is intentionally algorithm-facing: we view the estimation procedure that produced $q(d\theta)$ as exogenous, and our goal is to compute stable, scalable portfolio rules given this uncertainty description. It is also a commitment model: the investor commits at time 0 to a single feedback map and does not update $q$ during trading. As a result, one must distinguish between (i) $\theta$-conditional (full-information) optimality conditions and objects that would be available if $\theta$ were observable (infeasible under latent $\theta$), and (ii) $q$-aggregated conditions that characterize optimality within the $\theta$-blind admissible class. Throughout, $\theta$-conditional objects are used only for offline diagnostics (e.g., heterogeneity inspection and infeasible upper bounds), whereas our algorithms target a single deployable $\theta$-blind rule; see Section 2.2.
@@ 2.1 Model and ex–ante objective in estimated diffusion markets
We work on a filtered probability space $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in[0,T]},\mathbb{P})$ supporting Brownian motions of appropriate dimension. Time is continuous and runs over a fixed finite horizon $[0,T]$.
Deterministic-parameter reference (classical CRRA Merton).
There is one risk-free asset (money market account) with price process $B$ satisfying
$\frac{dB_{t}}{B_{t}}=r\,dt,\hskip 18.49988ptB_{0}=1,$ (1)
where $r\in\mathbb{R}$ is a constant short rate. In the classical Merton model, the $d$ risky assets have prices $S_{t}=(S_{t}^{1},\dots,S_{t}^{d})^{\top}$ solving
$\frac{dS_{t}}{S_{t}}:=\bigg(\frac{dS_{t}^{1}}{S_{t}^{1}},\dots,\frac{dS_{t}^{d}}{S_{t}^{d}}\bigg)^{\top}=r\,\mathbf{1}\,dt+\mu\,dt+\Sigma^{1/2}dW_{t},\hskip 18.49988ptS_{0}\in(0,\infty)^{d},$ (2)
with constant excess returns $\mu\in\mathbb{R}^{d}$, volatility matrix $\Sigma^{1/2}\in\mathbb{R}^{d\times d}$, and a $d$-dimensional Brownian motion $W$. An investor with CRRA utility $U(x)=x^{1-\gamma}/(1-\gamma)$, $\gamma>0,\gamma\neq 1$, chooses a progressively measurable portfolio fraction $\pi_{t}\in\mathbb{R}^{d}$; the wealth process satisfies
$\frac{dX_{t}^{\pi}}{X_{t}^{\pi}}=\big(r+\pi_{t}^{\top}\lambda\big)dt+\pi_{t}^{\top}\Sigma^{1/2}dW_{t},\hskip 18.49988ptX_{0}^{\pi}=x>0,$ (3)
where $\lambda:=\mu$ is the vector of risk premia. In this benchmark setting the optimal policy is constant:
$\pi^{\star}=\frac{1}{\gamma}\,\Sigma^{-1}\lambda,$ (4)
and the corresponding value function is given explicitly by
$V^{\mathrm{Merton}}(t,x;\lambda)=\frac{x^{1-\gamma}}{1-\gamma}\,\exp\Big\{(1-\gamma)\Big(r+\tfrac{1}{2\gamma}\lambda^{\top}\Sigma^{-1}\lambda\Big)(T-t)\Big\},$ (5)
see, for example, merton1969lifetime; merton1971optimum. We use this constant-coefficient model only as a deterministic-parameter reference.
@@
In our main formulation, drift and volatility are not assumed known. Instead, we consider a general multi-asset, multi-factor diffusion family indexed by $\theta\in\Theta\subset\mathbb{R}^{k}$, where $\theta$ represents the (possibly high-dimensional) parameter produced by an estimation procedure. Conditional on $\theta$, the $d$ risky assets and an $m$-dimensional factor process $Y_{t}$ evolve as
@@
The parameter $\theta$ is estimated from finite samples and is uncertain. We summarize this uncertainty by a probability distribution
@@
We deliberately do not tie $q$ to any specific inference paradigm. Concretely, $q$ may represent an empirical/sampling distribution produced by resampling procedures such as the bootstrap (efron1979bootstrap; efron1994bootstrap), a distribution induced by model averaging or sub-sample aggregation procedures such as bagging (breiman1996bagging), an approximate Bayesian posterior (when a prior and likelihood/criterion are specified), or an asymptotic normal (or sandwich) approximation in parametric or semiparametric estimation. For our purposes, $q$ is an input object describing statistically plausible market parameters.
@@
We interpret $\theta$ as a latent (unobserved) market parameter: Nature draws an $\mathcal{F}_{0}$-measurable random variable $\theta\sim q$ at time 0 (independent of the Brownian drivers) and keeps it fixed over $[0,T]$. The investor knows $q$ but does not observe the realized $\theta$, so deployable portfolio rules cannot take $\theta$ as an input.
@@
where $\mathcal{A}^{\mathrm{obs}}$ is defined below. This restriction reflects a fixed-$q$ commitment model: the investor uses historical data to form $q$ prior to trading and does not perform online filtering/belief-state updates during $[0,T]$.
@@
The $\theta$-blind constraint makes (14) strictly harder than solving a separate control problem for each fixed $\theta$, since the latter yields a $\theta$-indexed full-information family. Ex–ante averaging in (13) can also create gradient cancellation across heterogeneous parameter draws when one attempts to learn a single global policy end-to-end. While an $\mathcal{F}_{t}^{\mathrm{obs}}$-adapted policy could, in principle, filter $\theta$ online and solve a belief-state control problem (see, e.g., bensoussan1985optimal; pham2017dynamic), we do not pursue that formulation here.
@@
This subsection records the Hamiltonian structure underlying our projection step and clarifies what “Pontryagin first-order conditions” mean when the market parameter $\theta$ is latent and admissible controls are $\theta$-blind. In particular, we distinguish between (i) $\theta$-conditional (full-information) criticality conditions that would apply if $\theta$ were observable (and are therefore infeasible under latent $\theta$), and (ii) $q$-aggregated criticality conditions that characterize stationarity within the $\theta$-blind admissible class for the fixed-$q$ ex–ante objective. Our discussion follows standard stochastic control/PMP arguments for diffusion control (e.g. yong1999stochastic; fleming2006controlled; pham2009continuous). We also comment on the relationship to partial-information (belief-state) PMP, but we do not develop that formulation here.

### v2

@@ Abstract
We study continuous-time CRRA portfolio choice in diffusion markets with uncertain estimated coefficients. Nature draws a latent parameter from a given distribution and keeps it fixed; the investor cannot observe this parameter and must commit to a parameter-blind policy maximizing an ex-ante objective. We treat the uncertainty distribution as an inference-agnostic sampling input.
We develop a simulation-only two-stage solver. Stage 1 extends Pontryagin-Guided Direct Policy Optimization (PG-DPO) by sampling parameters internally and computing gradients via backpropagation through time. Stage 2 performs an aggregated Pontryagin projection: it aggregates costates across the parameter distribution to enforce a deployable stationarity condition, yielding a structured correction amortized via interactive distillation.
We prove a uniform conditional BPTT-PMP correspondence and a residual-based policy-gap bound with explicit error terms. Experiments on high-dimensional Gaussian drift and factor-driven benchmarks show that projection stabilizes learning and accurately recovers analytic references, while a model-free PPO baseline remains far from the targets.
Keywords: Portfolio choice; Parameter uncertainty; Deep learning; Pontryagin principle; Stochastic control
@@ 1 Introduction
A central problem in quantitative finance is to allocate wealth dynamically across many risky assets in continuous time. In the classical Merton model, investment opportunities are described by a low-dimensional diffusion with known drift and volatility, and closed-form optimal portfolios follow from HJB analysis; see, for example, Merton (1969, 1971). In realistic markets, however, key inputs—especially expected returns—must be estimated from finite samples, often in high dimensions and under regularization/model selection. Empirically, return predictability is unstable and many predictors deliver limited out-of-sample gains (e.g., Goyal and Welch, 2008; Campbell and Thompson, 2008). These considerations motivate separating diffusion risk (Brownian noise conditional on parameters) from decision-time estimation error (uncertainty in estimated coefficients), a theme that also underlies intertemporal hedging and more conservative allocations in models with predictability/learning (e.g., Kandel and Stambaugh, 1996; Barberis, 2000; Campbell and Viceira, 2002; Xia, 2001; Maenhout, 2004).
@@
We study continuous-time portfolio choice when market dynamics are known only up to an estimated parameter $\theta\in\Theta$, and an external estimation pipeline produces a nondegenerate uncertainty law $q(d\theta)$ over $\Theta$. We treat $q$ as an input object that summarizes decision-time estimation error about $\theta$. Concretely, $q$ may be obtained from resampling-based approximations (e.g., the bootstrap) (e.g., Efron, 1979; Efron and Tibshirani, 1994), from subsample aggregation and model-selection variability (e.g., bagging), or from asymptotic (sandwich/normal) approximations of an estimator’s sampling distribution (van der Vaart, 1998). Our goal is not to revisit inference, but to optimize decisions given this uncertainty description. Algorithmically, we interact with $q$ only through sampling $\theta\sim q$ inside a simulator; we do not assume a closed-form density, conjugacy, or a particular parametric family. We seek portfolio policies that maximize terminal CRRA utility ex–ante, averaging over both diffusion noise and parameter draws $\theta\sim q$.
@@
A key modeling choice is that $\theta$ is latent: Nature draws a fixed but unobserved $\theta\sim q$ at time 0 (independent of the Brownian drivers) and keeps it fixed on $[0,T]$. The investor knows $q$ but does not observe the realized $\theta$ and must therefore deploy a single $\theta$-blind policy. We restrict attention to Markov feedback policies of the form $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ that depend only on observable wealth $X_{t}$ and market factors $Y_{t}$, and we do not augment the state by a filtering/belief process. This is a one-shot decision-time problem: $q$ is treated as fixed over the trading horizon, and we do not model or hedge subsequent updates of the uncertainty description. Belief-state control (optimal learning/hedging of parameter uncertainty) is conceptually important, but it defines a different and typically far more demanding problem than computing a single deployable $\theta$-blind rule from a fixed $q$ (e.g., Bensoussan and van Schuppen, 1985; Pham and Wei, 2017).
@@
The $\theta$-blind deployability constraint also changes what a first-order optimality condition means. Under full information, Pontryagin’s Maximum Principle (PMP) yields a $\theta$-conditional criticality condition, hence a $\theta$-conditional feedback map that is infeasible when $\theta$ is latent. Under $\theta$-blind deployability, admissible perturbations are also $\theta$-blind, and the correct necessary condition is $q$-aggregated: the expectation over $\theta\sim q$ of the Hamiltonian gradient $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ must vanish along the controlled state process (e.g., Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009). In our portfolio setting, $\partial_{\pi}H^{\mathrm{ctrl}}_{\theta}$ is affine in $\pi$, so $q$-aggregation yields a statewise linear system whose solution defines a deployable $\theta$-blind projection map. This condition is inference-agnostic: it depends on $q$ only through its role as the ex–ante mixing law.
These features place the problem beyond the reach of classical dynamic programming in the targeted high-dimensional regime. Even in moderate dimensions, HJB solvers face the curse of dimensionality (e.g., Bellman, 1961; Kushner and Dupuis, 2001). Although deep PDE (e.g., Raissi et al., 2019; Sirignano and Spiliopoulos, 2018) and BSDE methods (e.g., Han et al., 2018; Beck et al., 2019) reduce grid reliance, nonlinear portfolio HJBs remain numerically delicate. Furthermore, belief augmentation renders the problem infinite-dimensional (e.g., Bensoussan and van Schuppen, 1985; Pham and Wei, 2017). Accordingly, we exclude value-function solvers as practical baselines under our $\theta$-blind interface, validating instead on closed-form benchmarks and simulation-only model-free RL (e.g., PPO).
@@
Our approach is simulation-based and builds on Pontryagin–Guided Direct Policy Optimization (PG–DPO) (Huh et al., 2025b, a). PG–DPO parameterizes a $\theta$-blind feedback policy via a neural network, simulates trajectories of the controlled SDE, and uses backpropagation through time (BPTT) to compute exact discrete-time gradients of terminal utility. Pathwise sensitivities computed by BPTT coincide with stochastic costates in PMP, reflecting the classical adjoint/backpropagation correspondence (e.g., LeCun, 1988; Yong and Zhou, 1999). In the latent-parameter setting, we sample $\theta\sim q$ inside the simulator (fixing it along each trajectory), while the policy depends only on observable states. To stabilize learning under heterogeneous $\theta$ draws, we develop uncertainty-aware projected P–PGDPO: after a warm-up phase that stabilizes costate estimates, we project Monte Carlo Pontryagin objects onto the $q$-aggregated PMP first-order condition, producing a robust deployable $\theta$-blind rule that can be amortized into a fast policy.
Empirically, in high-dimensional static drift experiments, the projected pipeline improves decision-time accuracy over end-to-end learning, stabilizing aligned regimes. In factor-driven markets where correlation induces hedging demand, it recovers analytic references under the $\theta$-blind restriction, whereas a PPO baseline remains far off. Theoretically, we establish a residual-based policy-gap bound: under mild stability, a small aggregated residual implies the policy is close to a locally optimal deployable rule, up to discretization and Monte Carlo errors.
@@
Our contributions are threefold. (i) We formulate a latent-parameter, fixed-$q$ ex–ante CRRA portfolio problem under a deployable $\theta$-blind Markov feedback restriction and derive the corresponding $q$-aggregated PMP first-order condition, emphasizing an inference-agnostic interface where uncertainty enters only through an exogenous mixing law $q(d\theta)$. (ii) We extend PG–DPO to this setting by sampling $\theta$ only inside the simulator and using BPTT to compute exact discrete-time gradients and pathwise sensitivities, and we establish a conditional BPTT–PMP correspondence uniform over $\theta$ on compact subsets of $\Theta$. (iii) We develop uncertainty-aware P–PGDPO that projects Monte Carlo costate estimates to produce a deployable $q$-aggregated $\theta$-blind rule, together with a residual-based ex–ante $\theta$-blind policy-gap bound and empirical evidence of two-time-scale stabilization.
@@
The remainder of the paper is organized as follows. Section 2 formulates the fixed-$q$ ex–ante problem and derives $\theta$-conditional versus $q$-aggregated PMP conditions together with Gaussian decision-time reference models. Section 3 develops PG–DPO and P–PGDPO for latent $\theta$ and proves the policy-gap bound. Sections 4 and 5 report numerical experiments, and the appendix collects technical proofs and implementation details.
@@ 2 Dynamic Portfolio Choice in Estimated Diffusion Markets with Latent Parameter Uncertainty
In this section we formalize a continuous-time CRRA portfolio problem in an estimated diffusion market with latent parameter uncertainty. Rather than fixing an inference architecture, we take as input an exogenous decision-time law $q(d\theta)$ over $\Theta$ and work with a latent, time-constant parameter $\theta\sim q$ (see Section 2.1 for the full model and objective). We adopt a $\theta$-blind commitment viewpoint: the investor commits at the decision time to a single Markov feedback rule $\pi_{t}=\bar{\pi}(t,X_{t},Y_{t})$ and does not update $q$ during trading. This restriction shifts the relevant optimality notion from infeasible $\theta$-conditional (full-information) conditions to a $q$-aggregated Pontryagin stationarity condition within the $\theta$-blind admissible class (Section 2.2). Closed-form Gaussian decision-time reference models used for validation are collected in Section 2.3.
@@ 2.1 Model and ex–ante objective in estimated diffusion markets
We interpret time 0 as the decision time and study portfolio choice on a fixed horizon $[0,T]$. The uncertainty law $q(d\theta)$ is an exogenous input available at the decision time and is treated as fixed over $[0,T]$ (no online belief updates during trading).
@@
Drift and volatility are not assumed known. Instead, we consider a general multi-asset, multi-factor diffusion family indexed by $\theta\in\Theta\subset\mathbb{R}^{k}$, where $\theta$ represents the (possibly high-dimensional) parameter produced by an estimation procedure. Conditional on $\theta$, the $d$ risky assets and an $m$-dimensional factor process $Y_{t}$ evolve as
@@
We summarize the uncertainty of the parameter $\theta$, estimated from finite samples, by a probability distribution
@@
We treat $q$ as inference-agnostic; it may represent resampling-based distributions (e.g., bootstrap (Efron, 1979; Efron and Tibshirani, 1994), bagging (Breiman, 1996)) or asymptotic approximations (van der Vaart, 1998). For our purposes, $q$ is a fixed decision-time input describing plausible market parameters over $[0,T]$.
@@
We interpret $\theta$ as a latent (unobserved) market parameter: at the (shifted) decision time 0, Nature draws an $\mathcal{F}_{0}$-measurable random variable $\theta\sim q$ (independent of the Brownian drivers) and keeps it fixed over $[0,T]$. The investor knows $q$ but does not observe the realized $\theta$, so deployable portfolio rules cannot take $\theta$ as an input.
@@
where $\mathcal{A}^{\mathrm{obs}}$ is defined below. This restriction reflects a fixed-$q$ commitment model: the investor forms $q$ at the decision time and does not perform online filtering/belief-state updates during $[0,T]$.
@@
The $\theta$-blind constraint makes (9) strictly harder than solving a separate control problem for each fixed $\theta$, since the latter yields a $\theta$-indexed full-information family. Ex–ante averaging in (8) can also create gradient cancellation across heterogeneous parameter draws when one attempts to learn a single global policy end-to-end. While an $\mathcal{F}_{t}^{\mathrm{obs}}$-adapted policy could, in principle, filter $\theta$ online and solve a belief-state control problem (see, e.g., Bensoussan and van Schuppen (1985); Pham and Wei (2017)), we do not pursue that formulation here.
@@
Decision-time commitment and (optional) plug-in replanning.
Throughout, we treat $q(d\theta)$ as an exogenous decision-time input and keep it fixed over the trading horizon $[0,T]$: within a single run, we do not update $q$ using newly observed returns or additional data, and we do not hedge against future changes of the uncertainty description. If the decision is revisited at a later calendar time $t_{0}\in(0,T)$, we interpret this as starting a new decision-time problem on the remaining horizon with an externally supplied law $q_{t_{0}}$.
Such a law may coincide with a static input (Section 2.3.1), may arise as a model-implied predictive (prior) propagation of an initial uncertainty through factor dynamics without conditioning on new observations (Section 2.3.2), or may be produced by an external filtering/estimation routine that outputs a time-indexed uncertainty description (Appendix A). Our model and algorithms are conditional on whichever $q$ is supplied at the decision time and always freeze that input over the ensuing horizon.
@@
This subsection records the Hamiltonian structure underlying our projection step and clarifies what “Pontryagin first-order conditions” mean when the market parameter $\theta$ is latent and admissible controls are $\theta$-blind. Throughout, $q(d\theta)$ denotes the decision-time input law (Section 2.1) and is treated as fixed over the trading horizon $[0,T]$. All expectations in this subsection are taken conditional on the decision-time information (suppressed in notation).
In particular, we distinguish between (i) $\theta$-conditional (full-information) criticality conditions that would apply if $\theta$ were observable (and are therefore infeasible under latent $\theta$), and (ii) $q$-aggregated criticality conditions that characterize stationarity within the $\theta$-blind admissible class for the fixed-$q$ ex–ante objective. Our discussion follows standard stochastic control/PMP arguments for diffusion control (e.g. Yong and Zhou, 1999; Fleming and Soner, 2006; Pham, 2009). We also comment on the relationship to partial-information (belief-state) PMP, but we do not develop that formulation here.